# Adaptive Prompt Elicitation for Text-to-Image Generation

### v1

@@ Abstract.
Aligning text-to-image generation with user intent remains challenging, for users who provide ambiguous inputs and struggle with model idiosyncrasies. We propose Adaptive Prompt Elicitation (APE), a technique that adaptively asks visual queries to help users refine prompts without extensive writing. Our technical contribution is a formulation of interactive intent inference under an information-theoretic framework. APE represents latent intent as interpretable feature requirements using language model priors, adaptively generates visual queries, and compiles elicited requirements into effective prompts. Evaluation on IDEA-Bench and DesignBench shows that APE achieves stronger alignment with improved efficiency. A user study with challenging user-defined tasks demonstrates 19.8% higher alignment without workload overhead. Our work contributes a principled approach to prompting that, for general users, offers an effective and efficient complement to the prevailing prompt-based interaction paradigm with text-to-image models.
@@ 1. INTRODUCTION
Our complementary evaluations provide convergent evidence for APE’s effectiveness and efficiency. On two text-to-image generation benchmarks, IDEA-Bench (Liang et al., 2025) and DesignBench (Hahn et al., 2025), APE achieves stronger alignment between generated images and ground-truth intents compared to baseline prompting approaches, while requiring significantly fewer interaction steps. A user study with 128 crowdworkers with varying experience levels further validates these findings in ecologically valid settings with self-defined creative tasks. Participants reported a 19.8% increase in perceived alignment between generated outputs and their intention compared to manual prompting. Qualitative analysis reveals that APE’s interactive elicitation process helps users discover and articulate preferences they could not initially verbalize, suggesting that the approach not only elicits existing intent but facilitates intent formation through guided exploration (Louie et al., 2020; Rezwana and Maher, 2023).
@@ 2.1. Prompt Engineering for Text-to-Image Generation
Complementing this line of work, automatic prompt optimization techniques have been proposed to help users write effective prompts without extensive learning. A prominent approach involves training specialized language models as prompt optimizers. Promptist (Hao et al., 2023) and BeautifulPrompt (Cao et al., 2023) combine supervised learning on curated preference data with reinforcement learning from AI feedback to improve semantic alignment and aesthetic quality. Prompt Expansion (Datta et al., 2024) focuses on enriching prompts with specialized keywords and additional details to enhance the aesthetic quality and diversity of generated images. More recently, large language models have been explored as black-box prompt optimizers (Liu et al., 2024; He et al., 2025; Mañas et al., 2024), leveraging their natural language understanding capabilities to refine user inputs.
@@ 2.2. Interaction Techniques in Text-to-Image Generation
Interaction techniques have emerged as a critical middle ground between manual prompt engineering and fully automatic optimization. By integrating human feedback into the generation loop, these systems aim to bridge the gap between abstract user intent and concrete model outputs. We categorize prior work into three primary paradigms: interface-driven exploration, mixed-initiative control, and query-based ambiguity resolution.
@@ 5.2.2. Qualitative Feedback
Participants’ primary complaints in the Baseline condition centered on the lack of alignment between their intent and the generated outputs, as well as the difficulty of achieving high-quality, fine-grained designs. Many reported that the system failed to fully interpret their prompts even after repeated refinements: “I keep adding [] to the prompt, but the generated image is not getting better at understanding the prompt.” Participants expressed frustration that the Baseline could not support detailed or precise adjustments: “I was only about 65 percent satisfied because even though it made the image I was going for, it wouldn’t let me tweak it further to take it to perfection.” Additionally, users frequently observed inconsistent adherence to instructions—such as incorrect numbers of people, unstable elements across iterations, or ignored constraints—which further undermined trust in the generation process: “It felt like some of the prompts I wrote were ignored. Thus the image I had in mind seemed to be impossible to be created.
@@ 5.2.3. The Effect of Prior Experience
To examine whether prior experience modulated the effect of APE on alignment ratings, we fitted an ordinal logistic regression model with condition (Baseline vs. APE), prior experience (modeled as an ordinal variable based on self-reported usage frequency), and their interaction. The model was fitted using maximum likelihood estimation ($N=128$, AIC = 917.9, BIC = 1023). The ordinal logistic regression revealed a significant main effect of condition ($\beta=2.26$, $SE=0.81$, $z=2.80$, $p=.005$), with participants using APE reporting higher alignment ratings than those in the Baseline condition. Prior experience was positively associated with alignment ratings, although this effect did not reach statistical significance ($\beta=0.32$, $SE=0.18$, $z=1.75$, $p=.080$). The interaction between condition and experience was not significant ($\beta=-0.29$, $SE=0.27$, z=1.08z=-1.08, $p=.280$), providing no evidence that the beneficial effect of APE on perceived alignment differed significantly across experience levels.
@@ 5.4.1. Interaction Patterns
Across iterations, users most commonly engage in repeated cycles of responding to visual outputs by selecting answer option, followed by a transition to image generation. While answer-to-answer transitions dominate early iterations, the relative proportion of transitions leading to generation increases over time, indicating that users tend to move from exploratory visual queries to generation after several rounds of refinement. Detours into manual requirement editing are comparatively infrequent. This pattern suggests that APE’s visual query interactions help users iteratively refine intent before committing to generation, reducing reliance on repeated text-based prompt rewriting.
@@ 5.4.3. Iteration Requirements
However, this advantage may be partially offset by increased latency due to additional API calls and background computation in APE. Our study employed FLUX.1, a diffusion-based model for rapid inference. One Baseline participant with prior experience using slower systems (Midjourney, DALL·E) commented: “I was impressed with how fast the images were generated,” suggesting FLUX’s speed may diminish the perceived value of iteration reduction. Conversely, APE’s additional computational overhead drew occasional criticism: “The system could be a little faster.” The efficiency trade-off likely shifts considerably with slower base models. For text-to-image systems requiring 30-60 seconds per generation, reducing iterations from 8 to 5 would save 1.5-3 minutes of cumulative wait time—a more substantial practical benefit. Future implementations could employ hybrid strategies: using fast distilled models (e.g., FLUX.1-schnell) for generating visual query exemplars during elicitation, then applying the refined prompt to slower, higher-quality models for final generation. This approach would preserve APE’s interaction efficiency while accommodating diverse model speed-quality trade-offs.
@@ 6. DISCUSSION
Across controlled technical evaluation (Section 4) and ecological user studies (Section 5), a consistent pattern emerges: information-theoretically optimized visual queries enable users to specify intent with greater precision than traditional prompt engineering, while requiring fewer iterations and imposing no additional cognitive burden. These benefits stem not from bypassing linguistic articulation, but from scaffolding it—transforming the task from open-ended prompt construction to guided refinement through comparative judgment. We now examine: (1) why visual query interaction proves effective for preference elicitation, (2) how transparent representations enhance user agency, (3) the tension between convergent elicitation and exploratory creativity, and (4) limitations and future directions.
@@ 6.1. Visual Query Interaction for Intent Elicitation
The reduction in image generation iterations represents substantial practical value in real-world use cases, consistent with prior work showing interactive query-based prompt optimization reduces search costs in image generation (Hahn et al., 2025; Wang et al., 2025). A generation cycle involves cognitive effort (evaluating results, deciding on modifications), time (waiting for generation), and computational cost (running the text-to-image model). By nearly halving the required iteration, APE transforms user experience from exhausting trial-and-error to structured refinement. This efficiency improvement appears particularly valuable for users with clear conceptual goals but limited prompt engineering strategies, though our exploratory subgroup analysis was underpowered for definitive claims about experience-based moderation.
@@ 6.4. Limitations and Future Work
Lastly, although APE is designed to be model-agnostic, our empirical evaluation is limited to a single state-of-the-art text-to-image model. Future work should examine its generalizability across a broader range of generative architectures to validate robustness and identify model-specific factors. In addition, investigating APE’s potential to extend across modalities (e.g., text-to-music and text-to-video generation) represents a promising direction for future research. As generative models continue to advance, addressing the alignment gap becomes increasingly critical beyond improvements in model capacity alone. APE constitutes an initial step toward mitigating this gap through principled interaction design, pointing toward future creative AI systems that are more interactive, transparent, and adaptive than current text-based interfaces.

### v2

@@ Abstract.
Aligning text-to-image generation with user intent remains challenging, as users frequently provide ambiguous inputs and struggle with model idiosyncrasies. We propose Adaptive Prompt Elicitation (APE), a technique that adaptively poses visual queries to help users refine prompts without extensive writing. Our technical contribution is a formulation of interactive intent inference under an information-theoretic framework. APE represents latent user intent as interpretable feature requirements using language model priors, adaptively generates visual queries, and compiles elicited requirements into effective prompts. Evaluation on IDEA-Bench and DesignBench shows that APE achieves stronger alignment with improved efficiency. A user study with 128 participants on user-defined tasks demonstrates 19.8% higher perceived alignment without increased workload. Our work contributes a principled approach to prompting that offers an effective and efficient complement to the prevailing prompt-based interaction paradigm with text-to-image models.
@@ 1. Introduction
Our complementary evaluations provide convergent evidence for APE’s effectiveness and efficiency. On two text-to-image generation benchmarks, IDEA-Bench (Liang et al., 2025) and DesignBench (Hahn et al., 2025), APE achieves stronger alignment between generated images and ground-truth intents compared to baseline prompting approaches, while requiring significantly fewer interaction steps. A user study with 128 participants with varying experience levels further validates these findings in ecologically valid settings with self-defined creative tasks. Participants reported a 19.8% increase in perceived alignment between generated outputs and their intention compared to manual prompting. Qualitative analysis reveals that APE’s interactive elicitation process helps users discover and articulate preferences they could not initially verbalize, suggesting that the approach not only elicits existing intent but facilitates intent formation through guided exploration (Louie et al., 2020; Rezwana and Maher, 2023).
@@ 2.1. Prompt Engineering for Text-to-Image Generation
Complementing this line of work, automatic prompt optimization techniques have been proposed to help users write effective prompts without extensive learning. A prominent approach involves training specialized language models as prompt optimizers. Promptist (Hao et al., 2023) and BeautifulPrompt (Cao et al., 2023) combine supervised learning on curated preference data with reinforcement learning from AI feedback to improve semantic alignment and aesthetic quality. Prompt Expansion (Datta et al., 2024) and UF-FGTG (Hei et al., 2024) train deep generative models to transform input prompts into high-quality, detailed descriptions using large-scale prompt expansion datasets, thereby enhancing the aesthetic quality and diversity of generated images. More recently, large language models have been explored as black-box prompt optimizers (Liu et al., 2024; He et al., 2025; Mañas et al., 2024; Chen et al., 2024), leveraging their natural language understanding capabilities to refine user inputs.
@@ 2.2. Interaction Techniques in Text-to-Image Generation
Interaction techniques have emerged as a critical middle ground between manual prompt engineering and fully automatic optimization. By integrating human feedback into the generation loop, these techniques seek to bridge the gap between abstract user intent and concrete model outputs. We categorize prior work into three primary paradigms: interface-driven exploration, mixed-initiative control, and query-based ambiguity resolution.
@@ 5.2.2. Qualitative Feedback
Participants’ primary complaints in the Baseline condition centered on the lack of alignment between their intent and the generated outputs, as well as the difficulty of achieving high-quality, fine-grained designs. Many reported that the system failed to fully interpret their prompts even after repeated refinements: “I keep adding [to the] prompt, but the generated image is not getting better at understanding the prompt.” Participants expressed frustration that the Baseline could not support detailed or precise adjustments: “I was only about 65 percent satisfied because even though it made the image I was going for, it wouldn’t let me tweak it further to take it to perfection.” Additionally, users frequently observed inconsistent adherence to instructions—such as incorrect numbers of people, unstable elements across iterations, or ignored constraints—which further undermined trust in the generation process: “It felt like some of the prompts I wrote were ignored. Thus the image I had in mind seemed to be impossible to be created.
@@ 5.2.3. The Effect of Prior Experience
To examine whether prior experience modulated the effect of APE on alignment ratings, we fitted an ordinal logistic regression model with condition (Baseline vs. APE), prior experience (modeled as an ordinal variable based on self-reported usage frequency), and their interaction. The model was fitted using maximum likelihood estimation ($N=128$, AIC = 917.9, BIC = 1023). The ordinal logistic regression revealed a significant main effect of condition ($\beta=2.26$, $SE=0.81$, $z=2.80$, $p=0.005$), with participants using APE reporting higher alignment ratings than those in the Baseline condition. Prior experience was positively associated with alignment ratings, although this effect did not reach statistical significance ($\beta=0.32$, $SE=0.18$, $z=1.75$, $p=.080$). The interaction between condition and experience was not significant ($\beta=-0.29$, $SE=0.27$, z=1.08z=-1.08, $p=.280$), providing no evidence that the beneficial effect of APE on perceived alignment differed significantly across experience levels.
@@ 5.4.1. Interaction Patterns
Across iterations, users most commonly engage in repeated cycles of responding to visual outputs by selecting an answer option, followed by a transition to image generation. While answer-to-answer transitions dominate early iterations, the relative proportion of transitions leading to generation increases over time, indicating that users tend to move from exploratory visual queries to generation after several rounds of refinement. Detours into manual requirement editing are comparatively infrequent. This pattern suggests that APE’s visual query interactions help users iteratively refine intent before committing to generation, reducing reliance on repeated text-based prompt rewriting.
@@ 5.4.3. Iteration Requirements
However, this advantage may be partially offset by increased latency due to additional API calls and background computation in APE. Our study employed FLUX.1, a flow-matching-based generative model for rapid inference. One Baseline participant with prior experience using slower systems (Midjourney, DALL·E) commented: “I was impressed with how fast the images were generated,” suggesting FLUX’s speed may diminish the perceived value of iteration reduction. Conversely, APE’s additional computational overhead drew occasional criticism: “The system could be a little faster.” The efficiency trade-off likely shifts considerably with slower base models. For text-to-image systems requiring 30-60 seconds per generation, reducing iterations from 8 to 5 would save 1.5-3 minutes of cumulative wait time—a more substantial practical benefit. Future implementations could employ hybrid strategies: using fast distilled models (e.g., FLUX.1-schnell) for generating visual query exemplars during elicitation, then applying the refined prompt to slower, higher-quality models for final generation. This approach would preserve APE’s interaction efficiency while accommodating diverse model speed-quality trade-offs.
@@ 6. Discussion
Across controlled technical evaluation (Section 4) and real-world user studies (Section 5), a consistent pattern emerges: information-theoretically optimized visual queries enable users to specify intent with greater precision than traditional prompt engineering, while requiring fewer iterations and imposing no additional cognitive burden. These benefits stem not from bypassing linguistic articulation, but from scaffolding it—transforming the task from open-ended prompt construction to guided refinement through comparative judgment. We now examine: (1) why visual query interaction proves effective for preference elicitation, (2) how transparent representations enhance user agency, (3) the tension between convergent elicitation and exploratory creativity, and (4) limitations and future directions.
@@ 6.1. Visual Query Interaction for Intent Elicitation
The reduction in image generation iterations represents substantial practical value in real-world use cases, consistent with prior work showing interactive query-based prompt optimization reduces search costs in image generation (Hahn et al., 2025; Wang et al., 2025). A generation cycle involves cognitive effort (evaluating results, deciding on modifications), time (waiting for generation), and computational cost (running the text-to-image model). By nearly halving the required iterations, APE transforms user experience from exhausting trial-and-error to structured refinement. This efficiency improvement appears particularly valuable for users with clear conceptual goals but limited prompt engineering strategies, though our exploratory subgroup analysis was underpowered for definitive claims about experience-based moderation.
@@ 6.4. Limitations and Future Work
Lastly, while the model-agnostic design of APE enables its application to a broad range of text-to-image generation models, our empirical evaluation is limited to a single state-of-the-art model. Future work could explore its generalizability across a wider set of generative architectures to assess robustness and uncover model-specific factors. In addition, investigating APE’s potential to extend across modalities (e.g. text-to-music (Copet et al., 2023) and text-to-video (Singer et al., 2023) generation) offers a compelling avenue for further study. As generative models continue to advance, addressing the alignment gap becomes increasingly critical beyond improvements in model capacity alone. APE constitutes an initial step toward mitigating this gap through principled interaction design, pointing toward future creative AI systems that are more interactive, transparent, and adaptive than current text-based interfaces.